if the difference of the two numbers is 9 and their product is 13 what is the sum of their squares/
Answer by mananth(16949)

if one number is x
the other number will be x+9
..
x(x+9)= 13
x^2 +9x =13
..
x^2 +(x+9)^2 is the sum of squares of the numbers
x^2 +x^2+18x+81
x^2 +9x +x^2 +9x +81
13+13+81= 107 = sum of squares.